OpenStudy (anonymous):

The following ordered pairs represents a function: (-2,10),(-1,7),(0,6),(1,7) and (2,10). What equation could represent the function?

OpenStudy (cwrw238):

Hint : (-1)^2 = 1 and (-2)^2 = 2^2

OpenStudy (cwrw238):

does that help?

OpenStudy (anonymous):

Not really for me sorry im not very good at math

OpenStudy (cwrw238):

square the first numbers in the ordered pairs and see how you can get the second numbers

OpenStudy (anonymous):

Ok i got it now thank you

OpenStudy (cwrw238):

yw